# von Neumann Orbit Equivalence

### v2

@@ 3.1. Von Neumann orbit equivalence for tracial von Neumann algebras
Since vNOE is an equivalence relation, it suffices to show that if $A\sim_{\mathrm{vNOE}}B$ and if $(C,\tau_{C})$ is another tracial von Neumann algebra, then $A*C\sim_{\mathrm{vNOE}}B*C$. Let $(Q,\tau_{Q})$ be a tracial von Neumann algebra and $\phi\colon B\to A\overline{\otimes}Q$ be a *-homomorphism as in condition (2) of Theorem 3.1. For tracial von Neumann algebra $(C,\tau_{C})$, we view $(A*C)\overline{\otimes}Q$ as $(A\overline{\otimes}Q)*_{Q}(C\overline{\otimes}Q)$. Define $\widetilde{\phi}_{0}\colon B*_{\textnormal{alg}}C\to(A\overline{\otimes}Q)*_{Q}(C\overline{\otimes}Q)$ by declaring that $\widetilde{\phi}_{0}(b)=\phi(b)*_{Q}1$ for $b\in B$ and $\widetilde{\phi}_{0}(c)=1*_{Q}c$ for $c\in C$. Let ${\mathbb{E}}_{Q}:(A*C)\,\overline{\otimes}\,Q\to Q$ be the normal conditional expectation. Note that ${\mathbb{E}}_{Q}\circ\phi=\tau_{B}$ and ${\mathbb{E}}_{Q}|_{C\,\overline{\otimes}\,Q}=\tau_{C}\otimes\operatorname{id}_{Q}$. Therefore, if $x\in B*_{\textnormal{alg}}C$ is an alternating centered word with respect to $\tau_{B*C}$, then $\widetilde{\phi}_{0}(x)$ is an alternating centered word with respect to ${\mathbb{E}}_{Q}$. Hence ${\mathbb{E}}_{Q}\circ\widetilde{\phi}_{0}=\tau_{B*C}$. Since ${\mathbb{E}}_{Q}$ is trace preserving, it follows that $\widetilde{\phi}_{0}$ is trace-preserving, and thus extends to a unique trace-preserving *-homomorphism $\widetilde{\phi}\colon B*C\to(A\overline{\otimes}Q)*_{Q}(C\overline{\otimes}Q)$. Moreover, by continuity we still have ${\mathbb{E}}_{Q}\circ\widetilde{\phi}=\tau_{B*C}$. In light of Remark 3.2, it thus remains to check that $\overline{\operatorname{Span}\{\widetilde{\phi}(x)y:x\in B*C,y\in Q\}}^{\|\cdot\|_{2}}=L^{2}((A*C)\,\overline{\otimes}\,Q)$. To this end, set $V=\overline{\operatorname{Span}\{\widetilde{\phi}(x)y:x\in B*C,y\in Q\}}^{\|\cdot\|_{2}}$. Since $V$ is invariant under right multiplciation by $Q$, to show that $V=L^{2}((A*C)\,\overline{\otimes}\,Q)$, it suffices to show that $(A*C)\otimes 1\subseteq V$. For this, by Lemma 3.6, it suffices to show that $V$ contains $A\otimes 1$ and $C\otimes 1$. That $C\otimes 1\subseteq V$, follows from the fact that $\widetilde{\phi}$ takes the copy of $C$ in $B*C$ to the copy of $C$ in $(A*C)\overline{\otimes}Q$, and since $\overline{\operatorname{Span}\{\phi(b)y:b\in B,y\in Q\}}^{\|\cdot\|_{2}}=L^{2}(A\overline{\otimes}Q)$, we also have that $A\otimes 1\subseteq V$. ∎
@@ 3.2. Von Neumann orbit equivalence for groups
Let $\Gamma{\,\curvearrowright\,}^{\sigma}\mathcal{M}$ be an action of a countable discrete group $\Gamma$ on a von Neumann algebra $\mathcal{M}$. A fundamental domain for the action is a projection $p\in\mathcal{M}$ such that the projections are $\{\sigma_{\gamma}(p)\}_{\gamma\in\Gamma}$ are pairwise orthogonal and $\sum_{\gamma\in\Gamma}\sigma_{\gamma}(p)=1$, where the sum converges in the strong operator topology. Two countable discrete groups $\Gamma$ and $\Lambda$ are said to be von Neumann equivalent, denoted $\Gamma\sim_{\mathrm{vNE}}\Lambda$, if there exists a semi-finite von Neumann algebra $(\mathcal{M},\mathrm{Tr})$ with a faithful normal semi-finite trace $\mathrm{Tr}$, and commuting trace-preserving actions $\Gamma{\,\curvearrowright\,}^{\sigma}\mathcal{M}$ and $\Lambda{\,\curvearrowright\,}^{\alpha}\mathcal{M}$ such that each action admits a finite-trace fundamental domain. Such an $\mathcal{M}$ is called a von Neumann coupling between $\Gamma$ and $\Lambda$.
@@
Two countable groups $\Gamma$ and $\Lambda$ are said to be von Neumann orbit equivalent, denoted $\Gamma\sim_{\mathrm{vNOE}}\Lambda$ if there exits a von Neumann coupling between $\Gamma$ and $\Lambda$ with a common fundamental domain.
@@
First suppose that $L\Gamma\sim_{\mathrm{vNOE}}L\Lambda$, and let $(\mathcal{H},Q,\xi)$ be a triple as in condition (1) of Theorem 3.1. Set $A=L\Gamma$ and $B=L\Lambda$, and consider $\mathcal{M}=Q^{\prime}\cap\mathcal{B}(H)={{}_{Q}\mathcal{B}(\mathcal{H})}$. For $\gamma\in\Gamma$, let $u_{\gamma}\in L\Gamma$ be the corresponding unitary and for $T\in\mathcal{B}(\mathcal{H})$, define $\sigma_{\gamma}(T)=u_{\gamma}Tu_{\gamma}^{*}$. Since $L\Gamma$- and $Q$-actions on $\mathcal{H}$ commute, it follows that $\mathcal{M}$ is invariant under $\sigma_{\gamma}$ for each $\gamma\in\Gamma$, and thus we have an action $\Gamma{\,\curvearrowright\,}^{\sigma}\mathcal{M}$. Similarly, since $\mathcal{H}$ is a $Q-B$-bimodule, we have an action $\Lambda{\,\curvearrowright\,}^{\alpha}\mathcal{M}$ given by $\alpha_{s}(T)=v_{s}^{*}Tv_{s},s\in\Lambda,T\in\mathcal{M}$, where $v_{s}\in L\Lambda$ is the unitary corresponding to $s\in\Lambda$. It is clear that the actions $\Gamma{\,\curvearrowright\,}^{\sigma}\mathcal{M}$ and $\Lambda{\,\curvearrowright\,}^{\alpha}\mathcal{M}$ commute. Let $\mathrm{Tr}$ be the canonical trace on $\mathcal{M}$ given by Proposition 2.4. It follows from the tracial property that both $\Gamma{\,\curvearrowright\,}^{\sigma}\mathcal{M}$ and $\Lambda{\,\curvearrowright\,}^{\alpha}\mathcal{M}$ are trace-preserving actions. Let $P\in\mathcal{B}(\mathcal{H})$ be the orthogonal projection from $\mathcal{H}$ onto $\overline{\operatorname{Span}(Q\xi)}$. It is straightforward to see that $P$ is $Q$-linear and thus, $P\in\mathcal{M}$. Moreover, it follows from Remark 2.5 that $\mathrm{Tr}(P)=1$. Therefore, it only remains to show that $P$ is a fundamental domain for both $\Gamma$- and $\Lambda$-actions. To see that $P$ is a $\Gamma$-fundamental domain, we first note that, since $\mathcal{H}=\overline{\operatorname{Span}(A\,\overline{\otimes}\,Q)\xi}$ and $\xi$ is tracial for the $A\,\overline{\otimes}\,Q$-module structure, we have, for $a\in A,x\in Q$, that
@@
If we let $P_{\gamma}$ be the orthogonal projection from $\mathcal{H}$ onto $u_{\gamma}(\overline{\operatorname{Span}(Q\xi)}$, then it follows from the above calculation that $\sigma_{\gamma}(P)=P_{\gamma}$, and it is straightforward to check that the projections $\{P_{\gamma}\}_{\gamma\in\Gamma}$ are pairwise orthogonal. Moreover, since $\mathcal{H}=\overline{\operatorname{Span}(A\,\overline{\otimes}\,Q)\xi)}$, we also get that $\sum_{\gamma\in\Gamma}\sigma_{\gamma}(P)=1$ and hence $P$ is a $\Gamma$-fundamental domain. Since we also have $\mathcal{H}=\overline{\operatorname{Span}(Q\xi B)}$ and $\xi$ is bi-tracial for the $Q-B$-bimodule structure, we observe that, for $b\in B,x\in Q$,
@@
If $b=\sum_{t\in\Lambda}b_{t}v_{t}$ is the Fourier series expansion of $b\in B$, then $\tau_{B}(b)=b_{e}$, and hence $P(x\xi b)=b_{e}x\xi$. For $s\in\Lambda$, let $P_{s}$ be the orthogonal projection from $\mathcal{H}$ onto $(\overline{\operatorname{Span}(Q\xi)})v_{s}$, then the following calculation show that $\alpha_{s}(P)=P_{s}$, $\{\alpha_{s}(p)\}_{s\in\Lambda}$ are pairwise orthogonal, and hence $P$ is a $\Lambda$-fundamental domain.
@@
$=(\alpha_{s}^{0}\otimes\lambda_{\Lambda}(s))(\overline{\operatorname{Span}\{y\sigma_{\gamma}(p)\otimes\delta_{e}:g\in\Gamma,y\in\mathcal{M}^{\Gamma}\}})$
@@
where $\rho_{\Lambda}:\Lambda\to\mathcal{U}(\ell^{2}\Lambda)$ is the right regular representation, makes $\mathcal{H}$ into a $Q-L\Lambda$-bimoudle. For $x\in\mathcal{M}^{\Gamma}$, and $s,t\in\Lambda$, we have
@@
and hence, $\xi$ is a tracial vector for the $Q-B$-bimodule structure. We recall from the proof of [IPR24, Proposition 4.2], that, since $p$ is $\Lambda$-fundamental domain, we have a direct sum decomposition $L^{2}(\mathcal{M},\mathrm{Tr})=\sum_{s\in\Lambda}L^{2}(\mathcal{M},\mathrm{Tr})\alpha_{s}(p)$. Thus, to show that $\overline{\operatorname{Span}(Q\xi B)}=\mathcal{H}$, it suffices to show that, for $s\in\Lambda$, $\overline{\operatorname{Span}\{xv_{s}\xi v_{t}:x\in\mathcal{M}^{\Gamma},t\in\Lambda\}}=L^{2}(\mathcal{M},\mathrm{Tr})\alpha_{s}(p)\,\overline{\otimes}\,\ell^{2}\Lambda$. To this end, fix $s\in\Lambda$ and note that

### v3

@@ 3.1. Von Neumann orbit equivalence for tracial von Neumann algebras
Since vNOE is an equivalence relation, it suffices to show that if $A\sim_{\mathrm{vNOE}}B$ and if $(C,\tau_{C})$ is another tracial von Neumann algebra, then $A*C\sim_{\mathrm{vNOE}}B*C$. Let $(Q,\tau_{Q})$ be a tracial von Neumann algebra and $\phi\colon B\to A\overline{\otimes}Q$ be a *-homomorphism as in condition (2) of Theorem 3.1. For tracial von Neumann algebra $(C,\tau_{C})$, we view $(A*C)\overline{\otimes}Q$ as $(A\overline{\otimes}Q)*_{Q}(C\overline{\otimes}Q)$. Define $\widetilde{\phi}_{0}\colon B*_{\textnormal{alg}}C\to(A\overline{\otimes}Q)*_{Q}(C\overline{\otimes}Q)$ by declaring that $\widetilde{\phi}_{0}(b)=\phi(b)*_{Q}1$ for $b\in B$ and $\widetilde{\phi}_{0}(c)=1*_{Q}c$ for $c\in C$. Let ${\mathbb{E}}_{Q}:(A*C)\,\overline{\otimes}\,Q\to Q$ be the normal conditional expectation. Note that ${\mathbb{E}}_{Q}\circ\phi=\tau_{B}$ and ${\mathbb{E}}_{Q}|_{C\,\overline{\otimes}\,Q}=\tau_{C}\otimes\operatorname{id}_{Q}$. Therefore, if $x\in B*_{\textnormal{alg}}C$ is an alternating centered word with respect to $\tau_{B*C}$, then $\widetilde{\phi}_{0}(x)$ is an alternating centered word with respect to ${\mathbb{E}}_{Q}$. Hence ${\mathbb{E}}_{Q}\circ\widetilde{\phi}_{0}=\tau_{B*C}$. Since ${\mathbb{E}}_{Q}$ is trace preserving, it follows that $\widetilde{\phi}_{0}$ is trace-preserving, and thus extends to a unique trace-preserving *-homomorphism $\widetilde{\phi}\colon B*C\to(A\overline{\otimes}Q)*_{Q}(C\overline{\otimes}Q)$. Moreover, by continuity we still have ${\mathbb{E}}_{Q}\circ\widetilde{\phi}=\tau_{B*C}$. In light of Remark 3.2, it thus remains to check that $\overline{\operatorname{Span}\{\widetilde{\phi}(x)y:x\in B*C,y\in Q\}}^{\|\cdot\|_{2}}=L^{2}((A*C)\,\overline{\otimes}\,Q)$. To this end, set $V=\overline{\operatorname{Span}\{\widetilde{\phi}(x)y:x\in B*C,y\in Q\}}^{\|\cdot\|_{2}}$. Since $V$ is invariant under right multiplication by $Q$, to show that $V=L^{2}((A*C)\,\overline{\otimes}\,Q)$, it suffices to show that $(A*C)\otimes 1\subseteq V$. For this, by Lemma 3.6, it suffices to show that $V$ contains $A\otimes 1$ and $C\otimes 1$. That $C\otimes 1\subseteq V$, follows from the fact that $\widetilde{\phi}$ takes the copy of $C$ in BCB*C to the copy of $C$ in $(A*C)\overline{\otimes}Q$, and since $\overline{\operatorname{Span}\{\phi(b)y:b\in B,y\in Q\}}^{\|\cdot\|_{2}}=L^{2}(A\overline{\otimes}Q)$, we also have that $A\otimes 1\subseteq V$. ∎
We conclude this subsection by proving Theorem 1.7, which we recall below.
Theorem 3.9.
Proof.
Let $(Q_{v},\phi_{v})$ be a vNOE-coupling between $(A_{v},\tau_{A_{v}})$ and $(B_{v},\tau_{B_{v}})$, where $\phi_{v}:A_{v}\to B_{v}\,\overline{\otimes}\,Q_{v}$ is a normal unital *-homomorphism satisfying
here, ${\mathbb{E}}_{v}:B_{v}\,\overline{\otimes}\,Q_{v}\to Q_{v}$ is the normal conditional expectation. Define $(Q,\tau_{Q})$ to be the tensor product $\overline{\otimes}\,_{v\in V}(Q_{v},\tau_{Q_{v}})$. We now define a *-homomorphism $\phi:A\to B\,\overline{\otimes}\,Q$ and show that it has the desired properties. To this end, first note that if $v$ and $w$ are two distinct vertices connected by an edge, then $B_{v}$ commutes with $B_{w}$ and $Q_{v}$ commutes with $Q_{w}$ as $v\neq w$. Thus, $\phi_{v}(A_{v})$ and $\phi_{w}(A_{w})$ commute inside $B\,\overline{\otimes}\,Q$, and therefore, we can define $\phi$ on the algebraic graph product (i.e., the universal unital *-algebra generated by the $\{A_{v}\}_{v\in V}$ subject to the relation that $A_{v}$ commutes with $A_{w}$ whenever $v$ and $w$ are connected by an edge). In particular, note that $\phi(a)=\phi_{v}(a)$, whenever $a\in A_{v}$. Next, we show that $\phi$ extends to a normal unital *-homomorphism on $A$ and satisfies ${\mathbb{E}}_{Q}(\phi(a))=\tau_{A}(a)$ for all $a\in A$, where ${\mathbb{E}}_{Q}:B\,\overline{\otimes}\,Q\to Q$ is the normal conditional expectation. To see this, we note that for any reduced word $v=v_{1}\cdots v_{n}$ of vertices, and elements $a_{i}\in A_{v_{i}}$ of trace zero, $1\leq i\leq n$, we have that
${\mathbb{E}}_{Q}(\phi_{v_{1}}(a_{1})\cdots\phi_{v_{n}}(a_{n}))=0.$
Indeed, for every $i\in\{1,\ldots,n\}$, we have ${\mathbb{E}}_{v_{i}}(\phi_{v_{i}}(a_{i}))=\tau_{A_{v_{i}}}(a_{i})=0,$ so, $\phi_{v_{i}}(a_{i})$ can be approximated by linear combinations of elements of the form $b_{v_{i}}\otimes q_{v_{i}}$, where $b_{v_{i}}\in B_{v_{i}},q_{v_{i}}\in Q_{v_{i}}$ with $\tau_{B_{v_{i}}}(b_{v_{i}})=0$. Thus, ${\mathbb{E}}_{Q}(\phi_{v_{1}}(a_{1})\cdots\phi_{v_{n}}(a_{n}))=0$ and hence it follows that the map $\phi$, defined on the algebraic graph product, is trace-preserving. Therefore, by [CF17, Proposition 3.22], $\phi$ extends to a well-defined normal unital *-homomorphism on $A$. Furthermore, if either $a=1$ or $a\in A$ is any reduced operator (see [CF17, Definition 3.10]), then we have seen that ${\mathbb{E}}_{Q}(\phi(a))=0=\tau_{A}(a)$. Since the linear span of $1$ and reduced operators is a dense *-subalgebra of $A$, we conclude that ${\mathbb{E}}_{Q}(\phi(a))=\tau_{A}(a)$ for all $a\in A$. Finally, to show that $\overline{\operatorname{Span}\{\phi(a)x:a\in A,x\in Q\}}^{\|\cdot\|_{2}}=L^{2}(B\,\overline{\otimes}\,Q)$, it suffices to show that
for any reduced word $v=v_{1}\cdots v_{n}$ of vertices and for all $n$. We proceed by induction. It is straightforward to see that for any vertex $v$, $\overline{\operatorname{Span}(\phi_{v}(A_{v})(1\otimes Q))}^{\|\cdot\|_{2}}=L^{2}(B_{v}\,\overline{\otimes}\,Q)=\overline{\operatorname{Span}(B_{v}(1\otimes Q))}^{\|\cdot\|_{2}}$. Suppose that (*3.1) holds for any reduced word $v_{1}\cdots v_{n}$ of length $n$ and let $v_{1}\cdots v_{n}v_{n+1}$ be any reduced word of length $n+1$. Then,
$\overline{\operatorname{Span}(\phi_{v_{1}}(A_{v_{1}})\cdots\phi_{v_{n+1}}(A_{n+1})(1\otimes Q))}^{\|\cdot\|_{2}}$ $=\overline{\operatorname{Span}(\phi_{v_{1}}(A_{v_{1}})B_{v_{2}}\cdots B_{v_{n+1}}(1\otimes Q))}^{\|\cdot\|_{2}}$
Thus, $(A,\tau_{A})\sim_{\mathrm{vNOE}}(B,\tau_{B})$. ∎
@@ 3.2. Von Neumann orbit equivalence for groups
Let $\Gamma{\,\curvearrowright\,}^{\sigma}\mathcal{M}$ be an action of a countable discrete group $\Gamma$ on a von Neumann algebra $\mathcal{M}$. A fundamental domain for the action is a projection $p\in\mathcal{M}$ such that the projections $\{\sigma_{\gamma}(p)\}_{\gamma\in\Gamma}$ are pairwise orthogonal and $\sum_{\gamma\in\Gamma}\sigma_{\gamma}(p)=1$, where the sum converges in the strong operator topology. Two countable discrete groups $\Gamma$ and $\Lambda$ are said to be von Neumann equivalent, denoted $\Gamma\sim_{\mathrm{vNE}}\Lambda$, if there exists a semi-finite von Neumann algebra $(\mathcal{M},\mathrm{Tr})$ with a faithful normal semi-finite trace $\mathrm{Tr}$, and commuting trace-preserving actions $\Gamma{\,\curvearrowright\,}^{\sigma}\mathcal{M}$ and $\Lambda{\,\curvearrowright\,}^{\alpha}\mathcal{M}$ such that each action admits a finite-trace fundamental domain. Such an $\mathcal{M}$ is called a von Neumann coupling between $\Gamma$ and $\Lambda$.
@@
Two countable groups $\Gamma$ and $\Lambda$ are said to be von Neumann orbit equivalent, denoted $\Gamma\sim_{\mathrm{vNOE}}\Lambda$ if there exists a von Neumann coupling between $\Gamma$ and $\Lambda$ with a common fundamental domain.
@@
First suppose that $L\Gamma\sim_{\mathrm{vNOE}}L\Lambda$, and let $(\mathcal{H},Q,\xi)$ be a triple as in condition (1) of Theorem 3.1. Set $A=L\Gamma$ and $B=L\Lambda$, and consider $\mathcal{M}=Q^{\prime}\cap\mathcal{B}(\mathcal{H})={{}_{Q}\mathcal{B}(\mathcal{H})}$. For $\gamma\in\Gamma$, let $u_{\gamma}\in L\Gamma$ be the corresponding unitary and for $T\in\mathcal{B}(\mathcal{H})$, define $\sigma_{\gamma}(T)=u_{\gamma}Tu_{\gamma}^{*}$. Since $L\Gamma$- and $Q$-actions on $\mathcal{H}$ commute, it follows that $\mathcal{M}$ is invariant under $\sigma_{\gamma}$ for each $\gamma\in\Gamma$, and thus we have an action $\Gamma{\,\curvearrowright\,}^{\sigma}\mathcal{M}$. Similarly, since $\mathcal{H}$ is a QBQ-B-bimodule, we have an action $\Lambda{\,\curvearrowright\,}^{\alpha}\mathcal{M}$ given by $\alpha_{s}(T)=v_{s}^{*}Tv_{s},s\in\Lambda,T\in\mathcal{M}$, where $v_{s}\in L\Lambda$ is the unitary corresponding to $s\in\Lambda$. It is clear that the actions $\Gamma{\,\curvearrowright\,}^{\sigma}\mathcal{M}$ and $\Lambda{\,\curvearrowright\,}^{\alpha}\mathcal{M}$ commute. Let $\mathrm{Tr}$ be the canonical trace on $\mathcal{M}$ given by Proposition 2.4. Since $L\Gamma$- and $Q$-actions on $\mathcal{H}$ commute, we have that for every $\gamma\in\Gamma$, $u_{\gamma}$ is a $Q$-linear operator on $\mathcal{H}$ and hence belongs to $\mathcal{M}$. Therefore, it now follows from the tracial property that $\mathrm{Tr}(\sigma_{\gamma}(T))=\mathrm{Tr}(u_{\gamma}Tu_{\gamma}^{*})=\mathrm{Tr}(T)$ for all $\gamma\in\Gamma,T\in\mathcal{M}$ and hence $\Gamma{\,\curvearrowright\,}^{\sigma}\mathcal{M}$ is trace-preserving. Similarly, $\Lambda{\,\curvearrowright\,}^{\alpha}\mathcal{M}$ too is trace-preserving.
Let $P\in\mathcal{B}(\mathcal{H})$ be the orthogonal projection from $\mathcal{H}$ onto $\overline{\operatorname{Span}(Q\xi)}$. It is straightforward to see that $P$ is $Q$-linear and thus, $P\in\mathcal{M}$. Moreover, it follows from Remark 2.5 that $\mathrm{Tr}(P)=1$. Therefore, it only remains to show that $P$ is a fundamental domain for both $\Gamma$- and $\Lambda$-actions. To see that $P$ is a $\Gamma$-fundamental domain, we first note that, since $\xi$ is tracial for the $A\,\overline{\otimes}\,Q$-module structure, we have, for $a\in A,x\in Q$, that
@@
If we let $P_{\gamma}$ be the orthogonal projection from $\mathcal{H}$ onto $u_{\gamma}(\overline{\operatorname{Span}(Q\xi)})$, then it follows from the above calculation that $\sigma_{\gamma}(P)=P_{\gamma}$, and it is straightforward to check that the projections $\{P_{\gamma}\}_{\gamma\in\Gamma}$ are pairwise orthogonal. Moreover, since $\mathcal{H}=\overline{\operatorname{Span}(A\,\overline{\otimes}\,Q)\xi}$, we also get that $\sum_{\gamma\in\Gamma}\sigma_{\gamma}(P)=1$ and hence $P$ is a $\Gamma$-fundamental domain. Since we also have that $\xi$ is bi-tracial for the QBQ-B-bimodule structure, we observe that, for $b\in B,x\in Q$,
@@
If $b=\sum_{t\in\Lambda}b_{t}v_{t}$ is the Fourier series expansion of $b\in B$, then $\tau_{B}(b)=b_{e}$, and hence $P(x\xi b)=b_{e}x\xi$. For $s\in\Lambda$, let $P_{s}$ be the orthogonal projection from $\mathcal{H}$ onto $(\overline{\operatorname{Span}(Q\xi)})v_{s}$. Now the following calculation shows that $\alpha_{s}(P)=P_{s}$, so, $\{\alpha_{s}(P)\}_{s\in\Lambda}$ are pairwise orthogonal, and thus $P$ is a $\Lambda$-fundamental domain since $\mathcal{H}=\overline{\operatorname{Span}(Q\xi B)}$.
@@
$=(\alpha_{s}^{0}\otimes\lambda_{\Lambda}(s))(\overline{\operatorname{Span}\{y\sigma_{g}(p)\otimes\delta_{e}:g\in\Gamma,y\in\mathcal{M}^{\Gamma}\}})$
@@
where $\rho_{\Lambda}:\Lambda\to\mathcal{U}(\ell^{2}\Lambda)$ is the right regular representation, makes $\mathcal{H}$ into a $Q-L\Lambda$-bimodule. For $x\in\mathcal{M}^{\Gamma}$, and $s,t\in\Lambda$, we have
@@
and hence, $\xi$ is a tracial vector for the QBQ-B-bimodule structure. We recall from the proof of [IPR24, Proposition 4.2], that, since $p$ is $\Lambda$-fundamental domain, we have a direct sum decomposition $L^{2}(\mathcal{M},\mathrm{Tr})=\sum_{s\in\Lambda}L^{2}(\mathcal{M},\mathrm{Tr})\alpha_{s}(p)$. Thus, to show that $\overline{\operatorname{Span}(Q\xi B)}=\mathcal{H}$, it suffices to show that, for $s\in\Lambda$, $\overline{\operatorname{Span}\{xv_{s}\xi v_{t}:x\in\mathcal{M}^{\Gamma},t\in\Lambda\}}=L^{2}(\mathcal{M},\mathrm{Tr})\alpha_{s}(p)\,\overline{\otimes}\,\ell^{2}\Lambda$. To this end, fix $s\in\Lambda$ and note that